what is an equation for a line perpendicular to y=-x-3 and passes through (0,0)?
Thank you :)

Answer by ewatrrr(24785)


Hi, Previously posted
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
equation for a line perpendicular to y=-x-3 and passes through (0,0)?
y = -x-3 ( m = (-1/1))
NEw Line: Perpendicular Line will have slope that is the negative reciprocal,
y = x + b |Using ordered pair (7,1) to solve for b
0 = 0 + b
0 = b
y = x